### v2

Project supported in part by the NSF of Shandong Province, China (No. ZR2019MA029),
the FRFCU(No. 3016000841964007), the NSF of Shandong Province, China (No.
ZR2014AM011) and the NSFC (No.11171184).
3. Proof of Theorem 9
First of all, we suppose, without loss of generality, that has finite order Combining this with Lemma 1, Lemma 2 (i) and the assumption that and are two distinct non-constant meromorphic functions that share IM and CM, we deduce that and are two distinct transcendental meromorphic functions in the complex plane.
From Lemma 2(i), Remark 4 and the assumption we have
From (16), Theorem 8 and Definition 1, we deduce that the orders of and are positive integers such that
Without loss of generality, we assume that and and set
By (17) and Whittaker [41, p. 82] we have
By (18), (19) and Lemma 4(i) we have
Since and share IM and CM, by simple calculating we can see that defined as (18) is analytic at any point such that for some and so from Lemma 3(i) and Remark 4 we have
From (20) and (21) we have
From (22) and Lemma 5 we see that is a rational function. Combining this with (18), we deduce that (18) can be rewritten into
where is reduced to zero or a non-vanishing polynomial, is a non-negative integer such that in the right hand side of (23) is reduced to zero when and with and is some non-zero integer when and while are those distinct points in such that and with and for some or and with and for some or with and is such a common zero of and in with different multiplicities for and but with and for all when and
By (18) and (23) we have
We integrate both sides of (24), and then we have
where is reduced to the constant when and when while is a complex constant or a non-constant polynomial such that
where is a finite complex constant. Next we rewrite (25) into
We consider the following two cases:
Case 1. Suppose that is reduced to the constant Then, it follows from (26) that is reduced to the finite complex constant Therefore, (27) can be rewritten into
On the other hand, by (17), Lemma 4 (i) and the second fundamental theorem (cf.[12, Theorem 2.1]) we have
Since and are transcendental meromorphic functions, by (29) and 5 we deduce that at least one of and say, satisfies
According to (30), Lemma 3 (ii) and the assumption that and share IM and CM, we have a discussion as follows:
Subcase 1.1. Suppose that
where and in what follows, denotes the reduced counting function of the common simple zeros of and in
By (31) we see that there exists
an a infinite sequence of common simple zeros of and in such that as Therefore, it follows by (28) that for each Combining this with the fact that is a non-vanishing rational function, we deduce that is reduced to the constant and so (28) can be rewritten into
By (32) and the assumption that and share IM we deduce that and share CM. This together with the assumption that and share CM, we get the assertion of Theorem 9.
Subcase 1.2. Suppose that there exists some positive integer satisfying such that
where and in what follows, denotes the reduced counting function of those common zeros of and in and each such point in is a simple zero of and a zero of of multiplicity By (33) we see that there exists an infinite sequence of those common zeros of and in such that as Here each such point is a simple zero of and a zero of of multiplicity Combining this with (28), we deduce
Since is a non-vanishing rational function, we deduce by (34) that is reduced to the constant and so (28) can be rewritten into
From (35), the assumption that and share IM, and the supposition that is a positive integer such that we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
and
are two entire functions. Therefore, from (17), (36) and Lemma 4 (i), we deduce
and
From (38) and (39) we deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that Then, it follows by (36) that
By (35) and (40) we deduce This contradicts the assumption that and are two distinct non-constant meromorphic functions.
Similarly, we get a contradiction from (35) and (37) provided that Next we suppose that and are non-vanishing polynomials. Integrating both sides of (36), we get
where is a finite complex constant. By (41) we have
Next we rewrite (37) into
Integrate both sides of (43) and we get
where is a finite complex constant. By (44) we have
From (35) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with for and Therefore, from (17), Lemma 2(ii), Remark 2, Lemma 4 (i) and the first fundamental theorem, we deduce
and so we have
Since and are transcendental meromorphic functions such that and share CM, we deduce by (16), (46) and Lemma 5 that
where and in what follows, denotes the reduced counting function of the common poles of and in
By (42) and (45) we deduce that if is a common pole of and then is a common zero of and Combining this with (47) and the assumption that and share CM, we deduce
From (42), (45) and the Valiron-Mokhonko lemma (cf.[23]) we deduce
and
as By (50) and (51) we have
By (16), (48) and (52) we have
By (49), (53) and Lemma 7, we see that there exist two integers and with such that This together with the definitions of and in (42) and (45) respectively gives
From (54) we have a discussion as follows:
Suppose that one of and is equal to zero, say Then Therefore, (54) can be rewritten into
We take the Nevanlinnas characteristic functions on both sides of (55), and then use the Valiron-Mokhonko lemma (cf.[23]) to get
By (16), (56) and the supposition that is a positive integer such that we have Combining this with
Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Suppose that Then Moreover, from we see that and are two non-zero inters. Therefore,
(54) can be rewritten into
We take the Nevanlinnas characteristic functions on both sides of (57) and then use the Valiron-Mokhonko lemma (cf.[23]) to get (56).
By (16), (56) and the supposition that is a positive integer such that we have Combining this with
Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Suppose that We first rewrite (54) into
We take the Nevanlinnas characteristic functions on both sides of (58) and then use the supposition and the Valiron-Mokhonko lemma (cf.[23]) to get
By (16), (59) and the supposition that is a positive integer such that we have Combining this with
Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Subcase 1.3. Suppose that there exists some positive integer satisfying such that
where and in what follows, denotes the reduced counting function of those common zeros of and in and each such point in is a zero of of multiplicity and a simple zero of By (60) we see that there exists an infinite sequence of those common zeros of and in such that as Here each such point is a zero of of multiplicity and a simple zero of Combining this with (28), we deduce
Since is a non-vanishing rational function, and is a complex constant, we deduce by (61) that is reduced to the constant and so (28) can be rewritten into
From (62), the assumption that and share IM, and the supposition that is a positive integer such that we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
and
are two entire functions. Next we use (17), (63), (64), Lemma 4 (i), the lines of (38) and (39) in Subcase 1.2 to deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that one of and holds. Then, we use (62)-(64) and the similar reasonings as in Subcase 1.2 to get a contradiction. Next we suppose that and are non-vanishing polynomials. Then, we use (63), (64), the the similar reasonings from the line before (41) to the line of (45) to deduce
and
where and in what follows, and are two finite and non-zero complex constants.
From (62) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with for and Therefore, from (17), Lemma 2(ii), Remark 2, Lemma 4 (i) and the first fundamental theorem, we deduce
and so we have
Since and are transcendental meromorphic functions such that and share CM, we deduce by (16), (67) and Lemma 5 that
By (65) and (66) we deduce that if is a common pole of and then is a common zero of and Combining this with (68) and the assumption that and share CM, we deduce
By (65), (66) and the Valiron-Mokhonko lemma (cf.[23]) we deduce
and
as By (71) and (72) we have
By (16), (69), (73), Lemma 5 and the obtained result that and are transcendental meromorphic functions we deduce
By (65), (66), (70), (74) and Lemma 7, we see that there exist some two integers and with such that This together with the definitions of and in (65) and (66) respectively gives
Next we use (75) and the similar reasonings from the line after (54) to the end of Subcase 1.2 to get a contradiction.
Case 2. Suppose that is a non-zero polynomial, then it follows by the definition of in (26) we see that is a non-constant polynomial. Next we let
where with are finite complex constants, and is a positive integer. Next we let with and and consider the following angles for any given positive number satisfying
where is an integer satisfying Then, it follows from (76) and Lemma 8 that there exists a positive number such that for we have
while
For convenience we next let
and
with
and
for each integer satisfying Then, we have the following claim:
Claim 2I. Based upon the assumptions of Theorem 9 and the supposition of Case 2, there are at most finitely many common zeros of and in for each integer and each integer satisfying
We prove Claim 2I: on the contrary, we suppose that Claim 2I is not valid. Then, there are infinitely many zeros of in for each integer and each integer satisfying Next we let be the infinite sequence
of all the zeros of on for and the integer satisfying where each distinct point in is repeated as many times as its multiplicity of a zero of with and the infinite sequence of is arranged according to increasing moduli, such that
and
for
On the other hand, since and are non-constant meromorphic functions of finite order such that they share IM and CM, we have by (17), Lemma 3(ii) and Remark 4 that
We consider the following two subcases:
Subcase 2.1. Suppose that with and is an even integer. Then, by (27), (77), (83), (84) and (85) we deduce that there exist some positive integer and a large positive number satisfying for and the even integer satisfying such that
and such that each point satisfying (86) is not a common multiple zero of and with and satisfies
as and for and the even integer satisfying Here and in what follows, denotes the multiplicity of of a zero of for the positive integer satisfying and the even integer satisfying Therefore, from (27), (85)-(87) and the assumption that and share and IM, we deduce that is also a simple zero of for the positive integer satisfying and the even integer satisfying Next we use Lemma 10 for to deduce
for the positive number such that From (86), (87) and Lemma 9 we have
for the the even integer satisfying and the positive number such that From (17) and Remark 5 we have
for the even integer satisfying and the positive number such that From (88)-(90) we deduce
for the even integer satisfying and the positive number such that
Next we use Lemma 9 and the reasonings in the proof of Lemma 2.2.2 in [49, pp.53-54] to deduce
for the even integer satisfying and the positive number such that Here and in what follows, for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (91) and (92) we have
for and the positive number such that
Now we use (86) and (87) to deduce for the positive integer satisfying Combining this with (83), we deduce that for any given positive number satisfying there exists some positive integer with that satisfies
such that for the positive integer satisfying with we have
From (95) we have
for the positive integer satisfying with From (93)-(96) we have
for the positive integer satisfying with From (83), (97) and the known equality with the even integer satisfying we have
for the even integer satisfying and the given positive number satisfying This is a contradiction.
Subcase 2.2. Suppose that with and is an odd integer. Then, by (27), (78), (83), (84) and (85) we deduce that there exist some positive integer and a large positive number satisfying for and the odd integer satisfying such that (86) holds and each point satisfying (86) is not a common multiple zero of and for and each point satisfying (86) satisfies
as for and the odd integer satisfying Here and in what follows, denotes the multiplicity of of a zero of for the positive integer satisfying and the odd integer satisfying From (27), (85), (99) and the assumption that and share and IM, we deduce that is also a simple zero of for the positive integer satisfying and the odd integer satisfying Next we use Lemma 10 for to deduce
for and the odd integer satisfying From (86), (99) and Lemma 9 we have
for the the odd integer satisfying and the positive number such that
From (17) and Remark 5 we have
for the odd integer satisfying and the positive number such that
From (100)-(102) we deduce
for the odd integer satisfying and the positive number such that Next we use Lemma 9 and the reasonings in the proof of
[49, pp.53-54, Lemma 2.2.2] to deduce
for the odd integer satisfying and the positive number such that Here and in what follows, for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (103) and (104) we have
for the odd integer satisfying and the positive number such that Now we use (86) and (99) to deduce for the positive integer satisfying Combining this with (83), we deduce that for the positive number satisfying there exists some large positive integer with that satisfies (94), such that for the positive integer satisfying with we have (95). From (95) we have
for and the positive integer satisfying From (94), (95), (105), (106) we have
for the odd integer satisfying and the positive integer satisfying with From (83), (107) and the known equality with the odd integer satisfying we deduce (98) for the odd integer satisfying and the given positive number satisfying From (98) we get a contradiction. This proves Claim 2I.
Last, we use Lemma 12, Lemma 13 and Claim 2I to complete the proof of Theorem 9. For this purpose, we have a discussion as follows: suppose that Then, it follows from Lemma 2.2 in [13] that there exist some real constant and some set that has infinite linear measure such that for all Combining this with Theorem 7, we see that and share all four values
CM, and so Theorem 9 is valid. Next we suppose that Then, it follows from the known equality that
From (108) we see that there exists some positive number satisfying
such that
where and in what follows, is a positive constant such that
Next we define the following directional arc:
with
for each integer satisfying From (109)-(111) we have
From (110)-(112) we see that there exists some integer satisfying say such that
Next we let
and let be the positive boundary of the simply connected domain where and are two oriented segments that are defined as
and
respectively, and is a directional arc that is defined as
where and in what follows, and are defined as in (82) for
Next we define as
Here and in what follows, is a small positive number satisfying such that Noting that with is continuous uniformly for on the closed interval and continuous at the origin point we deduce from (110) and (111) with (113) and (118) that
for the point the small positive number satisfying the positive number satisfying and the large positive number satisfying Here and in what follows, is defined as in (110) for From (110) and (111) with (113)-(119), the definitions of the proximity function and the Nevanlinna characteristic function of with the center at for we have
for the small positive number satisfying the positive number satisfying and the large positive number satisfying On the other hand, from Lemma 12, Claim 2I and the definition of the counting function of with the center at for we have
where
with and for and From (122), Claim 2I and Lemma 13 we see that there exists a set
that has linear measure zero, such that if then there is a constant, say that depend only upon and such that for all satisfying and and for each we have
From the formula (2.1.1) in [49, p.27], we have
From (123), (124) and the definitions of the proximity functions and the Nevanlinna characteristic functions of and for with the center at for we have
for with all satisfying
From (122) and (125) we deduce
for with all satisfying
From (121) and (126) we have
for with all satisfying
From (120) and (127) we have
for with all satisfying
From (128), Lemma 5, the supposition that is a positive number satisfying and the known result that and are transcendental meromorphic functions, we get a contradiction. This completes the proof of Theorem 9.