### v7

Project supported in part by the NSF of Shandong Province, China (No. ZR2019MA029),
the FRFCU(No. 3016000841964007), the NSF of Shandong Province, China (No.
ZR2014AM011) and the NSFC (No.11171184).
3. Proof of Theorem 7
First, before the formal beginning of the proof of Theorem 7, we want to point out that the lines from the next formal beginning of the proof of Theorem 7 to the end of Case 1 of the next formal beginning of the proof of Theorem 7 are covered essentially by the proof of Steinmetz [46, Theorem 6.2] even without any assumption on the order of growth. For the convenience of the readers, we give the detailed proof for this part on the assumption of finite order of growth as follows: first of all, we suppose, without loss of generality, that has finite order Combining this with Lemma 1, Lemma 2 (i) and the assumption that and are two distinct non-constant meromorphic functions that share IM and CM, we deduce that and are two distinct transcendental meromorphic functions in the complex plane. From Lemma 2(i), Remark 4 and the assumption we have
From (19), Theorem 6 and Definition 1, we deduce that the orders of and are positive integers such that
Without loss of generality, we assume that and and set
By (20) and Whittaker [50, p. 82] we have
By (21), (22) and Lemma 4 we have
Since and share IM and CM, by simple calculating we can see that defined as (21) is analytic at any point such that for some and so from Lemma 3(i) and Remark 4 we have
From (23) and (24) we have
From (25) and Lemma 5 we see that is a rational function. Combining this with (21), we deduce that (21) can be rewritten into
where is reduced to zero or a non-vanishing polynomial, is a non-negative integer such that in the right hand side of (26) is reduced to zero when and with and is some non-zero integer when and while are those distinct points in such that and with and for some or and with and for some or with and is such a common zero of and in with different multiplicities for and but with and for all when and
By (21) and (26) we have
We integrate both sides of (27), and then we have
where is reduced to the constant when and when while is a complex constant or a non-constant polynomial such that
where is a finite complex constant. Next we rewrite (28) into
On the other hand, from (20), Lemma 2(i), Lemma 3(ii), Lemma 4, Remark 2-Remark 4 and the assumption that and are non-constant meromorphic functions that share IM and CM, we have
from we have
and
Here and in what follows, denotes the counting function of the multiple poles of in From (31), (32) and Lemma 6 we have
as From (30) we consider the following two cases:
Case 1. Suppose that is reduced to the constant Then, it follows from (29) that is reduced to the finite complex constant Therefore, (30) can be rewritten into
According to (30), (33) and the assumption that and share IM and CM, we have a discussion as follows:
Subcase 1.1. Suppose that
Here and in what follows, denotes the reduced counting function of the common simple zeros of and in for the value From (35) we suppose, without loss of generality, that
From (36) we see that there exists
an a infinite sequence of the common simple zeros of and in the complex plane, such that as Then, it follows by (34) that for each Combining this with the fact that is a non-vanishing rational function, we deduce that is reduced to the constant and so (34) can be rewritten into
From (37) and the assumption that and share IM we deduce that and share CM. This together with the assumption that and share CM, we get the assertion of Theorem 7.
Subcase 1.2. Suppose that there exists some positive integer satisfying such that
where and in what follows, denotes the reduced counting function of those common zeros of and in that are simple zeros of and are zeros of of multiplicity for From (38) we suppose, without loss of generality, that
From (39) we see that there exists an infinite sequence of those common zeros of and in the complex plane that are simple zeros of and are zeros of of multiplicity Combining this with (34), we deduce
Since is a non-vanishing rational function, we deduce from (40) that is reduced to the constant and so (34) can be rewritten into
From (41), the assumption that and share IM, and the supposition that is a positive integer satisfying we deduce that if is a zero of of multiplicity for and then is a zero of of multiplicity for and This implies that the functions
and
are two entire functions. Therefore, from (20), (42) and Lemma 4, we deduce
and
From (44) and (45) we deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that Then, it follows by (42) that
By (41) and (46) we deduce This contradicts the assumption that and are two distinct non-constant meromorphic functions.
Similarly, we get a contradiction from (41) and (43) provided that Next we suppose that and are non-vanishing polynomials. Integrating both sides of (42), we get
where is a finite complex constant. By (47) we have
Next we rewrite (43) into
Integrate both sides of (49) and we get
where is a finite complex constant. By (50) we have
From (41) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with and for and Therefore, from (20), Lemma 2(ii), Remark 2, Lemma 4 and the first fundamental theorem, we deduce
and so we have
Since and are transcendental meromorphic functions such that and share CM, we deduce by (19), (52) and Lemma 5 that
where and in what follows, denotes the reduced counting function of the common poles of and in
By (48) and (51) we deduce that if is a common pole of and then is a common zero of and Combining this with (53) and the assumption that and share CM, we deduce
From (48), (51) and the Valiron-Mokhon′ko lemma (cf.[32]) we deduce
and
as By (56) and (57) we have
By (19), (54) and (58) we have
By (55), (59) and Lemma 7, we see that there exist two integers and with such that This together with the definitions of and in (48) and (51) respectively gives
From (60) we have a discussion as follows:
Suppose that one of and is equal to zero, say Then Therefore, (60) can be rewritten into
We take the Nevanlinna′s characteristic functions on both sides of (61), and then use the Valiron-Mokhon′ko lemma (cf.[32]) to get
By (19), (62) and the supposition that is a positive integer such that we have Combining this with
Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Suppose that Then Moreover, from we see that and are two non-zero inters. Therefore,
(60) can be rewritten into
We take the Nevanlinna′s characteristic functions on both sides of (63) and then use the Valiron-Mokhon′ko lemma (cf.[32]) to get (62).
By (19), (62) and the supposition that is a positive integer such that we have Combining this with
Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Suppose that We first rewrite (60) into
We take the Nevanlinna′s characteristic functions on both sides of (64) and then use the supposition and the Valiron-Mokhon′ko lemma (cf.[32]) to get
By (19), (65) and the supposition that is a positive integer such that we have Combining this with
Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Subcase 1.3. Suppose that there exists some positive integer satisfying such that
where and in what follows, denotes the reduced counting function of those common zeros of and in that are zeros of of multiplicity and are simple zeros of for From (66) we suppose, without loss of generality, that
From (67) we see that there exists an infinite sequence of those common zeros of and in the complex plane
that are zeros of of multiplicity and simple zeros of Combining this with (34), we deduce
Since is a non-vanishing rational function, and is a complex constant, we deduce from (68) that is reduced to the constant and so (34) can be rewritten into
From (69), the assumption that and share IM, and the supposition that is a positive integer satisfying we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
and
are two entire functions. Next we use (20), (70), (71), Lemma 4, the lines of (44) and (45) in Subcase 1.2 to deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that one of and holds. Then, we use (69)-(71) and the similar reasonings as in Subcase 1.2 to get a contradiction. Next we suppose that and are non-vanishing polynomials. Then, we use (70), (71), the the similar reasonings from the line before (47) to the line of (51) to deduce
and
where and in what follows, and are two finite and non-zero complex constants.
From (69) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with and for and Therefore, from (20), Lemma 2(ii), Remark 2, Lemma 4 and the first fundamental theorem, we deduce
and so we have
Since and are transcendental meromorphic functions such that and share CM, we deduce by (19), (74) and Lemma 5 that
From (72) and (73) we deduce that if is a common pole of and then is a common zero of and Combining this with (75) and the assumption that and share CM, we deduce
By (72), (73) and the Valiron-Mokhon′ko lemma (cf.[32]) we deduce
and
as By (78) and (79) we have
By (19), (76), (80), Lemma 5 and the obtained result that and are transcendental meromorphic functions we deduce
By (72), (73), (77), (81) and Lemma 7, we see that there exist some two integers and with such that This together with the definitions of and in (72) and (73) respectively gives
Next we use (82) and the similar reasonings from the line after (60) to the end of Subcase 1.2 to get a contradiction.
Case 2. Suppose that is a non-zero polynomial, then it follows by the definition of in (29) we see that is a non-constant polynomial. Next we let
where is a positive integer, and with are finite complex constants. Next we let with and and consider the following angles for any given positive number satisfying
where is an integer satisfying Then, it follows from (83) and Lemma 8 that there exists a positive number such that for we have
while
For convenience we next let
and
with
and
for each integer satisfying Then, we have the following claim:
Claim 2I. Based upon the assumptions of Theorem 7 and the supposition of Case 2, there are at most finitely many common zeros of and in for each integer and each integer satisfying
We prove Claim 2I: on the contrary, we suppose that Claim 2I is not valid. Then, there are infinitely many zeros of in for each integer and each integer satisfying Next we let be the infinite sequence
of all the zeros of on for and the integer satisfying where each distinct point in is repeated as many times as its multiplicity of a zero of with and the infinite sequence of is arranged according to increasing moduli, such that
and
for We consider the following two subcases:
Subcase 2.1. Suppose that with and is an even integer. Then, by (30), (84), (90), (91) and (32) we deduce that there exist some positive integer and a large positive number satisfying for and the even integer satisfying such that
and such that each point satisfying (92) is not a common multiple zero of and with and satisfies
as and for and the even integer satisfying Here and in what follows, denotes the multiplicity of of a zero of for the positive integer satisfying and the even integer satisfying Therefore, from (30), (32)-(93) and the assumption that and share and IM, we deduce that is also a simple zero of for the positive integer satisfying and the even integer satisfying Next we use Lemma 10 for to deduce
for the positive number such that From (92), (93) and Lemma 9 we have
for the the even integer satisfying and the positive number such that From (20) and Lemma 12 we have
for the even integer satisfying and the positive number such that From (94)-(96) we deduce
for the even integer satisfying and the positive number such that
Next we use Lemma 9 and the reasonings in the proof of Zheng [57, pp.53-54, Lemma 2.2.2] to deduce
for the even integer satisfying and the positive number such that Here and in what follows,
for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (97) and (98) we have
for and the positive number such that
Now we use (92) and (93) to deduce
for the positive integer satisfying Combining this with (90), we deduce that for any given positive number satisfying there exists some positive integer with that satisfies
such that for the positive integer satisfying with we have
From (102) we have
for the positive integer satisfying with From (100)-(103) we have
for the positive integer satisfying with From (90), (104) and (99) with the even integer satisfying we have
for the even integer satisfying and the given positive number satisfying This is a contradiction.
Subcase 2.2. Suppose that with and is an odd integer. Then, by (30), (85), (90), (91) and (32) we deduce that there exist some positive integer and a large positive number satisfying for and the odd integer satisfying such that (92) holds and each point satisfying (92) is not a common multiple zero of and for and each point satisfying (92) satisfies
as for and the odd integer satisfying Here and in what follows, denotes the multiplicity of of a zero of for the positive integer satisfying and the odd integer satisfying From (30), (32), (106) and the assumption that and share and IM, we deduce that is also a simple zero of for the positive integer satisfying and the odd integer satisfying Next we use Lemma 10 for to deduce
for and the odd integer satisfying From (92), (106) and Lemma 9 we have
for the the odd integer satisfying and the positive number such that
From (20) and Lemma 12 we have
for the odd integer satisfying and the positive number such that
From (107)-(109) we deduce
for the odd integer satisfying and the positive number such that Next we use Lemma 9 and the reasonings in the proof of
Zheng [57, pp.53-54, Lemma 2.2.2] to have
for the odd integer satisfying and the positive number such that Here and in what follows, is defined as with being the number of zeros of for in counted with multiplicities. From (110) and (111) we have
for the odd integer satisfying and the positive number such that Now we use (92) and (106) to deduce for the positive integer satisfying Combining this with (90), we deduce that for the positive number satisfying there exists some large positive integer with that satisfies (101), such that for the positive integer satisfying with we have (102). From (102) we have
for and the positive integer satisfying From (101), (102), (112), (113) we have
for the odd integer satisfying and the positive integer satisfying with From (90), (114) and (99) with the odd integer satisfying we deduce (105) for the odd integer satisfying and the given positive number satisfying From (105) we get a contradiction. This proves Claim 2I.
Next we use Lemma 16, Lemma 18 and Claim 2I to complete the proof of Theorem 7. For this purpose, we have a discussion as follows: suppose that Then, it follows from Lemma 2.2 in [21] that there exist some real constant and some set that has infinite linear measure such that for all Combining this with Theorem 5, we see that and share all four values
CM, and so Theorem 7 is valid. Next we suppose that Then, it follows from the known equality that
Next we define the following directional arc:
with
for each integer satisfying From (116) and (117) we have
On the other hand, from (20), (115) and Lemma 15 we have
From (115)-(119) we see that there exists some integer satisfying say such that
with
for arbitrary Pó1ya
peaks of order as Here and in what follows, is some positive constant, and is some set of finite linear measure. Then, it follows from (120) and (121) that
where and in what follows, is a large positive number satisfying
Next we set the simply connected domain
with the positive boundary where and are two oriented segments such that
and
respectively, and is a directional arc such that
where and in what follows, and are defined as (89) for Next we also define with as
for Here and in what follows, is a small positive number satisfying and
is a complex number such that Then, from (123)-(128), Remark 5.3, the positive boundary of the region and the definition of the proximity function of with the center at for we have
for the small positive number satisfying Next we also set the following transformation
where and in what follows, is a simply connected domain such that
where is a positive number satisfying
and the positive boundary of the simply connected domain is
where and in what follows, is an oriented segment, and is a directional arc. They are defined as follows: the oriented segment is such an oriented segment that for each we have with and the starting point of is while the end point of is
The directional arc is such a directional arc that for each we have
with and the starting point of is with while the end point of is with In addition, from (89) with we have
From (130) and (133) we can easily verify that the transformation of (130) maps the simply connected domain of (131) onto the simply connected domain of (123) in a one-to-one mapping manner, and maps the positive boundary of into the positive boundary of in a one-to-one mapping manner,
where the oriented segment and the directional arc are mapped into and respectively by the transformation of (130) in a one-to-one mapping manner. Moreover, from (131)-(133), the polar form of the Cauchy-Riemann partial differential equations and the polar form of the sufficient conditions for an analytic function (cf.[43]), we can verify that the transformation of (130) is also an analytic transformation in the simply connected domain of (131). Combining this with the known result that the transformation of (130) maps the simply connected domain of (131) onto the simply connected domain of (123) in a one-to-one mapping manner, we see that the univalent transformation (130) on the simply connected domain is a univalent and analytic map on the simply connected domain and maps conformally the simply connected domain of (131) with the positive boundary onto the simply connected domain of (123) with the positive boundary For convinience of the following discussion, we denote
the univalent and analytic transformation (130) as
and there exists a unique point satisfying
such that the univalent and analytic mapping (134) maps conformally the simply connected domain of (131) with the positive boundary onto the simply connected domain of (123) with the positive boundary
On the other hand, from the Riemann mapping theorem (cf.[2, p.230,Theorem 1]) we also see that there exists a unique univalent and analytic function on the simply connected domain of (123), say
such that the univalent and analytic transformation (136) maps conformally the simply connected domain of (123) with the positive boundary onto the open unit disk
with the positive boundary with where the starting point of is with and the end point of is with From (134)-(136) we deduce that the composition
of the transformation (136) and the transformation (134) with for
is a univalent and analytic transformation of onto the open unit disk such that it maps the positive boundary of into the positive boundary of the open unit disk mentioned above in a one-to-one mapping manner. Noting that the transformation of (136) is a univalent and analytic mapping of onto the open unit disk Combining this with the theorem on the boundary behavior of conformal mappings, we see that the function
(137) is continuous on and for Next we use the similar reasoning of proving (8) in Remark 5 to have
and
for each point with and on the semi-circle
Since the transformation (134) is a univalent and analytic transformation of onto such that (135) holds, we have from (134), (136)-(138), (140) and Lemma 17 that the Green’s functions and for and with singularities and respectively satisfy the relation
for each
Next we use (135), (137)-(142) and the invariance of first order differential form to deduce
with for each point with and on the semi-circle
From (128), (131), (132), (141), (13) in Remark 5.1 and Goldberg-Ostrovskii [14, p.306] we have
for each point with and on the semi-circle
and the point
Here and in what follows
From (122), (129), (135), (143), (144)-(146), the supposition the definitions of the proximity function and the Nevanlinna characteristic function of with the center at for for and we deduce
i.e.,
for the small positive number satisfying and the positive number and Here and in what follows,
Next we denote by
the Nevanlinna′s angular characteristic function on the angular domain
with
Next we use (149), (150) and the line of the reasoning for proof of (97) in Subcase 2.1 of the proof of Claim 2I to deduce
for the positive number such that say. From (122), (149)-(151) and Lemma 13, we deduce that there exists some positive constant and there exists some set of finite logarithmic measure, such that when and the relation
holds uniformly in Here and in what follows
Similarly, from Lemma 13 and (97) for we deduce that there exists some positive constant and there exists some set of finite logarithmic measure, such that the relation
holds uniformly in when and Here and in what follows, is defined as (99) for From (128), (154) and the supposition we deduce
as and On the other hand, from (147), (150), (152) and (153) we have
as and Last, from Lemma 16 we have
where
with and for Noting that denotes the Green function for the half-disc
we use the formula of the Green function (12) for the half-disc
in Remark 5.1 to have
with
for each From (130), (142) and (159) we have
for each with
From Claim 2I we see that there exist finitely many distinct zeros of with in say
such that
where is some positive integer that depends only upon the positive numbers and Next we suppose that the multiplicity of of a zero of is equal to some positive integer for and Then, from (161), (146), the reasoning in Goldberg-Ostrovskii [14, p.306], the supposition and the definition (5) of the counting function for the meromorphic function with the center at for we have
as
Here and in what follows,
for and On the other hand, from Lemma 18 we see that there exists some set
of finite logarithmic measure such that
as and From the formula (2.1.1) in Zheng [57, p.27], we have
From (164), (165) and the definitions of the proximity functions and the Nevanlinna characteristic functions of and for with the center at for we have
as and From (128) and (155) we deduce
as and From Remark 4, Lemma 3 (i), the assumption and the assumption that and share IM and CM we have
which implies that and have finitely many critical values including and possibly in the complex plane. Combining this with Lemma 20 and the assumption we deduce that the set of finite singular values of the inverse function is bounded. Therefore, from Lemma 19 we see that there exist constants and such that if and then
where is a positive absolute constant, in particular independent of and From (128), (155), (168) and the supposition we see that there exists some set of finite logarithmic measure such that
as and From (133) and (99) for we have Combining this with (128), (155) and (169) we deduce
as and From (128), (155), (170), Lemma 18 and the supposition
we see that there exists some set of finite logarithmic measure, such that
as and From (158), (166), (167), (171) we have
as and From (157), (158), (162), (163), (172) we have
as and as
From (146), (156) and (173) we have
as and From (148), (153), (89) and (99) for we see that
for any given small positive number satisfying
From (174), (175) and (146) we derive a contradiction, when and This completes the proof of Theorem 7.