### v8

# On a question of Gary G. Gundersen concerning meromorphic functions sharing three distinct values IM and a fourth value CM

Abstract.
In 1992, Gundersen [18] proposed the following famous open question: if two non-constant meromorphic functions share three values IM and share a fourth value CM, then do the functions necessarily share all four values CM? The open question is a long-standing question in the studies of the Nevanlinna′s value distribution theory of meromorphic functions, and has not been completely resolved by now. In this paper, we prove the following result: suppose that and are two distinct non-constant meromorphic functions, and one of and has finite order. If and share IM and CM, where are four distinct complex values in the extended complex plane, then and share and CM. Applying the main result obtained in this paper, we resolve a question proposed in Gundersen [15, p.458] concerning the nonexistence of two distinct non-constant meromorphic functions sharing three distinct values DM and a fourth value CM for finite order meromorphic functions. The obtained result also extends the corresponding result in Mues[34, pp.109-117] concerning the nonexistence of two distinct and non-constant entire functions that share three distinct finite values DM. Examples are provided to show that the main results obtained in this paper, in a sense, are best possible.
1. Introduction and main results
In this paper, by meromorphic functions we will always mean meromorphic functions in the complex plane. We adopt the standard notations in the Nevanlinna theory of meromorphic functions as explained in [20, 23, 50]. Throughout this paper, we denote by a set of finite linear measure. For a meromorphic function in the complex plane we denote by any quantity satisfying as and Following Halburd-Korhonen [19, p.466], we define the notion of the small function of a non-constant meromorphic function in the complex plane as follows: let be a non-constant meromorphic function. Next we denote by the set of all meromorphic functions in the complex plane such that for all possibly outside a set of finite linear measure. Functions in the set are called small compared to or slowly moving with respect to or a small function of for short.
Let and be two non-constant meromorphic functions in the complex plane, and let be a value in the extended plane. We say that and share the value CM, provided that and have the same set of -points, and each common -point of and has the same multiplicities as the -point of and We say that and share the value IM, provided that and have the same set of -points ignoring multiplicities. We say that and share the value DM, provided that and share IM, and at least one common -point of and has different multiplicities related to and (cf.[15, 16]). Here we mention that Mues [34] and Yang-Yi [50] defined the notion of the DM shared value as follows: we say that and share the value DM, provided that and share IM, and that and have no common zeros of the same multiplicities. For a non-constant meromorphic function we denote by and respectively the lower order and the order of the relevant definitions can be found, for example, in [20, 23, 50]. For convenience, we recall them below:
Definition 1.
For a non-constant meromorphic function the lower order and the order of denoted as and respectively are defined as
Following Edrei-Fuchs [10, p.293] , we say that a meromorphic function in the complex plane is of regular growth, provided Here we mention that both of and may be
In 1926, Nevanlinna [37] proved the following theorems that are the famous five-value theorem and the four-value theorem respectively:
Theorem 1.
([37, p.109]) If and are non-constant meromorphic functions that share five values IM, then .
Theorem 2.
([37, p.122]) Let and be distinct and non-constant meromorphic functions, and let be four distinct complex values in the extended complex plane. If and share CM, then is a Möbius transformation of , two of the shared values, say and , are Picard exceptional values, and the cross ratio
In 1976, L. Rubel posed the following question(cf.[15]): whether do the four distinct CM shared values be replaced with four distinct IM shared values or not in Theorem 2? In this direction, Gundersen [15] and Gundersen [16] proved the following results respectively, where Theorem 4 improved Theorems 2 and 3, and Theorem 3 improved Theorem 2:
Theorem 3.
([15, Theorem 1]) If and share three values CM and share a fourth value IM, then they share all four values CM.
Theorem 4.
([16, Theorem 1]) If two non-constant meromorphic functions share two values CM and share two other values IM, then the functions share all four values CM.
Gundersen [15] gave the following famous example that shows that the L.Rubel’s question above is negative:
Example 1.
([15, pp.458-459]) Let and where is a non-constant entire function, we can easily verify that and share DM, and is not a Möbius transformation of
Based upon Example 1 and Theorem 4, Gundersen ([15, 16, 18]) proposed the following famous question:
Question 1.
Question 1 is the main open question in the theory of meromorphic functions that share four values. This open question appears to be difficult and has not been completely resolved by now. In the past several decades, there are many partial results on Question 1. For the ease of the following narrations, we introduce the following notation: let and be two non-constant meromorphic functions, and let be a value in the extended complex plane. Suppose that and share IM. Next we denote by the counting function of those common -points of and in where each point in has the same multiplicities regard to and and each point in is counted only once. We define
for the large positive number where if From the definition we see that We mention that is defined as when and share CM. In 1989, Mues [35, Theorem 1] proved that if and are two distinct and non-constant meromorphic functions in the complex plane that share four distinct values in the extended complex plane, and one of the four distinct values is shared by and CM, and for another value of the four distinct shared values, then and share all the four distinct values CM. In 1993, Wang [48] improved [35, Theorem 1] and proved that if and are two distinct and non-constant meromorphic functions in the complex plane that share four distinct values in the extended complex plane, and two of the four distinct values, say and satisfy and then and share all the four distinct values CM. As far as we know, there are other research works concerning Question 1 and the uniqueness question of meromorphic functions that share four distinct values in the complex plane, such as Mues [34, 36], Gundersen [18], Reinders [40], Ueda [46], Wang[47], Steinmetz [44, 45], Huang[25] and Huang-Du[26], etc. In addition, Ishizaki[22], Li [29] and Yao [53] studied the uniqueness question of non-constant meromorphic functions sharing four distinct small functions. We recall the following partial results in Gundersen [18, Theorem 1] and Li-Yi [30] on the open question respectively:
Theorem 5.
([18, Theorem 1]) Let and be two non-constant meromorphic functions that share IM and CM, where and are four distinct complex values in the extended complex plane. Suppose that there exist some real constant and some set that has infinite linear measure such that for all Then and share all four values CM. Here when
Theorem 6.
([18, Theorem 1]) Let and be two distinct and non-constant meromorphic functions, and let be four distinct values in the extended complex plane. If and share IM and CM, then and are functions of normal growth, and have the same order, and the order of and is a positive integer or infinite.
Next we consider the following special case of Question 1:
Question 2.
If two non-constant meromorphic functions share three values IM and share a fourth value CM, where one of the two non-constant meromorphic functions has finite order, then do the two functions necessarily share all four values CM?
In this paper, we give an affirmative answer to Question 2 for two distinct and non-constant meromorphic functions satisfying the assumptions in Question 2. Indeed, we prove the following theorem in this paper:
Theorem 7.
Suppose that and are two distinct and non-constant meromorphic functions, and one of and has finite order. If and share IM and CM, where are four distinct complex values in the extended complex plane, then and share and CM.
Remark 1.
From Theorem 7 we get the following result:
Corollary 1.
Suppose that and are two distinct and non-constant entire functions, and one of and has finite order. If and share IM, where are three distinct finite complex values, then and share and CM.
2. Preliminaries
In this section, we will introduce some results that play an important role in proving the main result in this paper. First of all, we introduce the following result due to Adams-Straus[1]:
Lemma 1.
([1, Theorem 3]). Let and be two non-constant rational functions. If and share four distinct values IM, then
The following result was proved in Nevanlinna [37] originally:
Lemma 2.
([37, p.373, Satz 3] or[15, Theorem 2]) Let and be two distinct and non-constant meromorphic functions, and let be three distinct and finite complex values in the complex plane. If and share and IM, then
(i) (ii)
Remark 2.
The following result and its proof can be found in Gundersen [16]:
Lemma 3.
([16, p.550, Lemma 2] and [16, p.549, Corollary 1(II)]). Let and be two diatinct and non-constant meromorphic functions that share four distinct values and IM, where Then the following statement holds:
(i) where and “count” respectively only those points in and which do not occur when for some
(ii) where with and denotes the counting function of those common -points of and in that are multiple for both and and each such point in is counted according to the number of the times of the smaller of the two multiplicities.
Remark 3.
A proof of Lemma 3 can be found in the proof of Lemma 3 in [18, p.102]. Following the expression of the Mues′s function in [13, pp.176-177] and the proof of Lemma 3 in [18, p.102], we can see that the error term in Lemma 3 is expressed as
where with is a finite complex constant that depends only on and and is called the E.Mues’s function that is an entire function defined as
when and share IM. This can be found in Mues [35, p.171, proof of Lemma 1]. Here we mention that based upon the assumption that two distinct and non-constant meromorphic functions and share and IM, where is a finite complex number such that the Mues′s function in Mues [35, p.171, proof of Lemma 1] is written into
The following result is due to Yang [52]:
Lemma 4.
([52, Theorem 1.6]). Suppose that is a non-constant meromorphic function of finite order, and let be a positive integer. Then
Remark 4.
The following result is from Yang-Yi[50]:
Lemma 5.
([50, p.11, Theorem 1.5]). If is a transcendental meromorphic function in the complex plane, then
Lemma 6.
([30, proof of Lemma 9]). Let and be two distinct and non-constant meromorphic functions, and one of and has finite order. Suppose that and share IM, where are three distinct finite values and Then
as Here and in what follows, denotes the reduced counting function of those common zeros of and in that are simple zeros of and are zeros of of multiplicity for with and with while denotes the reduced counting function of those common zeros of and in that are zeros of of multiplicity and are simple zeros of for and
Lemma 7.
([54, Lemma 6]). Let and be two non-constant meromorphic functions such that
Then, either or there exist two integers and with such that Here and in what follows, denotes the reduced counting function of the common -points in and is any quantity such that as and where and denotes a set of finite linear measure.
The following result is due to Markushevich [31]:
Lemma 8.
([31]). Let where is a positive integer and with and For any given positive number satisfying we consider angles:
where is an integer satisfying Then, there exists a positive number such that for with where is even, and for with where is odd.
Next we introduce the Nevanlinna theory of meromorphic functions in one angular domain that will play a key role in the proof of the main result: let be a meromorphic function on the angular domain where Following Goldberg-Ostrovskii [14, p.25], we define
where and with and is the sequence of poles of on and each distinct point in the sequence is repeated as many times as its multiplicity of a pole of on We denote by the reduced form of In other word, if each point in the above expression of is counted only once for each distinct point in the sequence we denote instead of We call the Nevanlinna′s angular counting function of poles of on The Nevanlinna′s angular characteristic function is defined as We recall the following two results are due to Zheng [55]:
Lemma 9.
([55, Lemma 3]). Let be meromorphic on Then for arbitrary complex number we have
Lemma 10.
([55, Lemma 4]). Let be meromorphic on Then for arbitrary distinct in we have
where the term will be replaced with when for some positive integer satisfying
From Zheng [56, p.81, Lemma 2.5.3] and the formula
in Zheng [56, p.56], we get the following result:
Lemma 11.
([56, Lemma 2.5.3]). Let be a meromorphic function in the complex plane. For any positive number such that we have
where with and is a constant independent of and
Lemma 12.
Based upon the assumptions of Lemma 11, we have the following conclusion: suppose that is a non-constant meromorphic function of finite order in the complex plane. Then, we have when and
Proof.
Suppose that is a non-constant meromorphic function of finite order in the complex plane. Then, from Lemma 4 we see that there exists a large positive number and a positive constant such that and when with where is a positive constant such that Combining this with Lemma 4, the formula (4) in Lemma 11 for and we have
as Here and in what follows, is a positive constant such that This reveals the conclusion of Lemma 12. ∎
The following result was originally proved in Goldberg-Ostrovskii[14]:
Lemma 13.
([14, p.291, Theorem 2.7]) Let be a meromorphic function on the angular domain where If then there exists a set of finite logarithmic measure, such that when and the relation
holds uniformly in with Here is a real constant and
Lemma 14.
([11, p.85], [51, Lemma 1] and [56, p.5, Theorem 1.1.3]) Let be a transcendental and meromorphic meromorphic function in the complex plane with the lower order and the order Then, for arbitrary positive number satisfying and a set with finite logarithmic measure, there exist a sequence of positive numbers such that
(1) and (2)
(3) for with
A sequence satisfying (I), (2) and (3) in Lemma 14 is called Pó1ya peaks of order outside with finite logarithmic measure. For and we define
and
The following result is a special version of the main result of Baernstein [3]:
Lemma 15.
([3, p.419]) Let be a transcendental and meromorphic function in the complex plane with the finite lower order and the order and for some point we have Then, for arbitrary Pó1ya peaks of order satisfying we have
Here and in what follows, is a set of finite logarithmic measure.
Next we introduce the Nevanlinna theory in a simply connected domain on (cf.[56, pp.26-38]): let be a simply connected domain surrounded by finitely many piecewise analytic curves. Then for any there exists a Green function, denoted as for with singularity at which is uniquely determined by the following conditions: (1) is harmonic in (2) in a neighborhood of for some function harmonic in (3) on the boundary of
Next we denote by the positive boundary of and the inner normal of with respect to Since for and for from the definition of directional derivative it follows that the directional derivative of on in the inner normal is non-negative, that is to say, with From the Green formula, in view of the Green function, we can establish the following formula, which is an extension of the Poisson formula for a disk. For a generalization of the formula, the reader is referred to [14, p.4, Theorem 1.1]. Following Zheng[56, p.28], we define the counting function of with the center at for as follows:
where is a point such that and with is a pole of appearing often according to its multiplicities of a pole of and is the multiplicity of a pole of at while is the sum in (5) counting all distinct with and with replaced by when Following Zheng [56, p.28], we also define
and
which are called the proximity function and the Nevanlinna characteristic function of with the center at for the simply connected domain respectively. For meromorphic functions on a simply connected domain surrounded by finitely many piecewise analytic curves, where is a positive integer such that the basic properties concerning the proximity functions, the Nevanlinna counting functions and the Nevanlinna characteristic functions of meromorphic functions on with the center at for the simply connected domain are similar to the corresponding properties concerning the classical proximity functions, the classical Nevanlinna counting functions and the classical Nevanlinna characteristic functions of meromorphic functions on the closed disc with the center at the origin point for the closed disk this can be found, for example, in Zheng [56, pp.28-29]. We recall the following result that is called the Nevanlinna second fundamental theorem with the center at for a simply connected domain (cf.[56, p.32]):
Lemma 16.
(cf.[56, p.32, Theorem 2.1.4]) Let be a meromorphic function on where is a simply connected domain such that and let be distinct finite complex numbers, where is a positive integer such that Then for such that and with and we have
where
and
Here and for with
We recall the following result that depicts the relationship between two Green’s functions for two domains of which is mapped to another one by a univalent analytic function:
Lemma 17.
(cf.[9, p.276, 5.3 Theorem]) Let and be regions on the complex plane such that there is a one-to-one analytic function of onto and let and If and are the Green’s functions for and with singularities and respectively, then for each
Remark 5.
Following Zheng [56, pp.35-37] and Goldberg-Ostrovskii[14, p.2], we have a discussion as follows: let be a non-constant meromorphic function on where is a simple connected domain that is surrounded by finitely many piecewise analytic curves, in other words, the boundary of the simple connected domain consists of finitely many piecewise analytic curves. Then, for a point such that by the Riemann mapping theorem (cf.[2, p.230,Theorem 1]) there exists a conformal mapping of onto the unit disc in such a way that By the theorem on the boundary behavior of conformal mappings, the function is continuous on and for It is easy for us to see that with satisfies (1) and (2). Along the positive boundary of we have
when is analytic in the boundary of Indeed, from the Cauchy-Riemann condition that the tangent vector at any point along the positive boundary of becomes the inner normal of at the corresponding point on after is rotated anticlockwise. Therefore, for each we have
when with and is an analytic function on Therefore, it follows from (9) and with that
for when the function mentioned above is analytic in the boundary of Here and in what follows, denotes the tangent vector along the positive boundary of at any given point and denotes the corresponding inner normal of the positive boundary of at the corresponding point From (10) we get (8). In this way, we can obtain the Green functions for some special simple connected domains on as follows:
(5.1) For and a point where is some positive number, we follow Zheng [56, p.36] and Goldberg-Ostrovskii[14, p.2, Example 2.] to have
and
for Moreover, from (8), (11), (12) and Goldberg-Ostrovskii[14, p.306] we have for each point with on the semi-circle and with and that
while for each point in the interval and with and we have
(5.2) For and the point where is any given positive number, we follow Zheng [56, p.36] to write the proximity function the Nevanlinna counting function the Nevanlinna characteristic function of with the center at for into and respectively. In particular, the proximity function the Nevanlinna counting function the Nevanlinna characteristic function of with the center at for are reduced into and respectively. Indeed, we substitute (11) into (8), and then we have
for any complex number with and any given complex number such that and a point such that In particular, we have from (14) that
for any complex number with Therefore, it follows from (5)-(7) and (15) that the proximity function the Nevanlinna counting function the Nevanlinna characteristic function of with the center at for are reduced into
and
respectively.
(5.3)Based upon the assumptions of Lemma 16, we follow Zheng [56, p.35] to see that and are non-negative for the case of while for the case of may be negative. For example, we consider the function with where is a positive integer. In view of (5), (16)-(18) and Theorem 2.16 in Zheng [56, p.34] we deduce and with for It is obvious that and so when for
For convenience in stating the following result, we shall use the following notation. We shall let denote a pair that consists of a transcendental meromorphic function and a finite set of distinct pairs of integers that satisfy for We recall the following result due to Gundersen [17]:
Lemma 18.
([17, p.90, Corollary 2]) Let be a given pair where has finite order and let be a given constant. Then, there exists a set that has finite logarithmic measure, such that if and for all we have
Let be a transcendental and meromorphic function in the complex plane. An important role in complex dynamics [4] is played by the singular values of the inverse function these are the critical values of and the asymptotic values, that is, values such that tends to as tends to infinity along a path We denote by the class of all transcendental and meromorphic functions in the complex plane for which the set of finite singular values of is bounded, and by we denote the subclass of consisting of those for which has finitely many singular values. Both classes and have been studied extensively in iteration theory and value distribution theory (cf.[4, 5, 12, 27, 41]). We recall the following result from Bergweiler [4] and Rippon [41] that was proved first in Eremenko [12] for entire functions:
Lemma 19.
The following result is due to Bergweiler-Eremenko [6]:
Lemma 20.
([6, p.90, Corollary 3]) If a meromorphic function of finite order has only finitely many critical values, then it has at most asymptotic values.
Next we introduce the notion of the spherical derivative of a meromorphic function in the complex plane (cf.[20, 52]): let be a non-constant meromorphic function. The spherical derivative of at is given as We recall the following result from Chang-Zalcman [7]:
Lemma 21.
([7, Lemma 1]). Let be a meromorphic function in If has bounded spherical derivative in then is of order at most If, in addition, is an entire function, then the order of is at most
Remark 6.
Following He-Xiao [24, pp.53-55], we introduce the Ahlfors-Shimizu’s characteristic function of a meromorphic function: suppose that is a meromorphic function in the complex plane. Then, the Ahlfors-Shimizu′s characteristic function of denoted as is defined as
where with The Nevanlinna’s characteristic function and the Ahlfors-Shimizu’s characteristic function differ by a bounded quantity that is independent of (cf.[24, p.56]). This implies the first part of Lemma 21. The result of Lemma 21 for entire functions is much subtler, which is a special case of Clunie-Hayman [8, Theorem 3].
We also need some notions of normal families of meromorphic functions in a domain of the complex plane. which can be found, for example in Montel [33] and Hayman [20, pp.157-160]. For the convenience of the readers, we give the detail of the relevant notions as follows: following Montel [33], we call a class of meromorphic functions in a domain to be normal in the domain provided that for any given sequence of meromorphic functions in we can find a subsequence which converges everywhere in and uniformly on compact subsets of with respect to the chordal metric on Riemann sphere. We then say that converges locally uniformly in An equivalent statement is that for every in there exists a neighbourhood in which or converges uniformly as This implies that if and are two points in the -plane, their distance in the chordal metric of Riemann sphere is
Also
Thus, if either or uniformly in a set then uniformly in Thus, if one of these two conditions holds uniformly in some neighbourhood of every point of then uniformly in some neighbourhood of every point of and hence by the Heine-Borel theorem uniformly on every compact subset of see, for example, Hayman [20, pp.157-160]. We recall the following result from Steinmetz [45]:
Lemma 22.
([45, Rescaling Lemma 7.1.]). Let and be two non-constant meromorphic functions in the extended complex plane, and let be four distinct complex values in the extended complex plane such that and share IM. Then, either the spherical derivatives and are uniformly bounded on or else there exist sequences and with as such that the sequences and with and being defined as and respectively for each simultaneously tend to non-constant meromorphic functions and in the complex plane respectively, such that and have uniformly bounded spherical derivatives on and such that and share the four complex values IM.
3. Proof of Theorem 7
First, before the formal beginning of the proof of Theorem 7, we want to point out that the lines from the next formal beginning of the proof of Theorem 7 to the end of Case 1 of the next formal beginning of the proof of Theorem 7 are covered essentially by the proof of Steinmetz [45, Theorem 6.2] even without any assumption on the order of growth. For the convenience of the readers, we give the detailed proof for this part on the assumption of finite order of growth as follows: first of all, we suppose, without loss of generality, that has finite order Combining this with Lemma 1, Lemma 2 (i) and the assumption that and are two distinct and non-constant meromorphic functions that share IM and CM, we deduce that and are two distinct transcendental meromorphic functions in the complex plane. From Lemma 2(i), Remark 4 and the assumption we have
From (19), Theorem 6 and Definition 1, we deduce that the orders of and are positive integers such that
Without loss of generality, we assume that and and set
By (20) and Whittaker [49, p. 82] we have
By (21), (22) and Lemma 4 we have
Since and share IM and CM, by simple calculating we can see that defined as (21) is analytic at any point such that for some and so from Lemma 3(i) and Remark 4 we have
From (25) and Lemma 5 we see that is a rational function. Combining this with (21), we deduce that (21) can be rewritten into
where is reduced to zero or a non-vanishing polynomial, is a non-negative integer such that in the right hand side of (26) is reduced to zero when and with and is some non-zero integer when and while are those distinct points in such that and with and for some or and with and for some or with and is such a common zero of and in with different multiplicities for and but with and for all when and By (21) and (26) we have
We integrate both sides of (27), and then we have
where is reduced to the constant when and when while is a complex constant or a non-constant polynomial such that
where is a finite complex constant. Next we rewrite (28) into
On the other hand, from (20), Lemma 2(i), Lemma 3(ii), Lemma 4, Remark 2-Remark 4 and the assumption that and are non-constant meromorphic functions that share IM and CM, we have from we have
and
Here and in what follows, denotes the counting function of the multiple poles of in From (31), (32) and Lemma 6 we have
as From (30) we consider the following two cases:
Case 1. Suppose that is reduced to the constant Then, it follows from (29) that is reduced to the finite complex constant Therefore, (30) can be rewritten into
According to (30), (33) and the assumption that and share IM and CM, we have a discussion as follows:
Subcase 1.1. Suppose that
Here and in what follows, denotes the reduced counting function of the common simple zeros of and in for the value From (35) we suppose, without loss of generality, that
From (36) we see that there exists an a infinite sequence of the common simple zeros of and in the complex plane, such that as Then, it follows by (34) that for each Combining this with the fact that is a non-vanishing rational function, we deduce that is reduced to the constant and so (34) can be rewritten into
From (37) and the assumption that and share IM we deduce that and share CM. This together with the assumption that and share CM, we get the assertion of Theorem 7.
Subcase 1.2. Suppose that there exists some positive integer satisfying such that
where and in what follows, denotes the reduced counting function of those common zeros of and in that are simple zeros of and are zeros of of multiplicity for From (38) we suppose, without loss of generality, that
From (39) we see that there exists an infinite sequence of those common zeros of and in the complex plane that are simple zeros of and are zeros of of multiplicity Combining this with (34), we deduce
Since is a non-vanishing rational function, we deduce from (40) that is reduced to the constant and so (34) can be rewritten into
From (41), the assumption that and share IM, and the supposition that is a positive integer satisfying we deduce that if is a zero of of multiplicity for and then is a zero of of multiplicity for and This implies that the functions
and
are two entire functions. Therefore, from (20), (42) and Lemma 4, we deduce
and
as From (44) and (45) we deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that Then, it follows by (42) that
By (41) and (46) we deduce This contradicts the assumption that and are two distinct and non-constant meromorphic functions. Similarly, we get a contradiction from (41) and (43) provided that Next we suppose that and are non-vanishing polynomials. Integrating both sides of (42), we get
where is a finite complex constant. By (47) we have
Next we rewrite (43) into
Integrate both sides of (49) and we get
where is a finite complex constant. By (50) we have
From (41) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with and for and Therefore, from (20), Lemma 2(ii), Remark 2, Lemma 4 and the first fundamental theorem, we deduce
and so we have
Since and are transcendental meromorphic functions such that and share CM, we deduce by (19), (52) and Lemma 5 that
where and in what follows, denotes the reduced counting function of the common poles of and in
By (48) and (51) we deduce that if is a common pole of and then is a common zero of and Combining this with (53) and the assumption that and share CM, we deduce
From (48), (51) and the Valiron-Mokhon′ko lemma (cf.[32]) we deduce
and
By (19), (54) and (58) we have
By (55), (59) and Lemma 7, we see that there exist two integers and with such that This together with the definitions of and in (48) and (51) respectively gives
From (60) we have a discussion as follows:
Suppose that one of and is equal to zero, say Then Therefore, (60) can be rewritten into
We take the Nevanlinna′s characteristic functions on both sides of (61), and then use the Valiron-Mokhon′ko lemma (cf.[32]) to get
By (19), (62) and the supposition that is a positive integer such that we have Combining this with Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Suppose that Then Moreover, from we see that and are two non-zero inters. Therefore, (60) can be rewritten into
We take the Nevanlinna′s characteristic functions on both sides of (63) and then use the Valiron-Mokhon′ko lemma (cf.[32]) to get (62). By (19), (62) and the supposition that is a positive integer such that we have Combining this with Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Suppose that We first rewrite (60) into
We take the Nevanlinna′s characteristic functions on both sides of (64) and then use the supposition and the Valiron-Mokhon′ko lemma (cf.[32]) to get
By (19), (65) and the supposition that is a positive integer such that we have Combining this with Lemma 5 and the supposition that is a transcendental meromorphic function in the complex plane, we get a contradiction.
Subcase 1.3. Suppose that there exists some positive integer satisfying such that
where and in what follows, denotes the reduced counting function of those common zeros of and in that are zeros of of multiplicity and are simple zeros of for From (66) we suppose, without loss of generality, that
From (67) we see that there exists an infinite sequence of those common zeros of and in the complex plane that are zeros of of multiplicity and simple zeros of Combining this with (34), we deduce
Since is a non-vanishing rational function, and is a complex constant, we deduce from (68) that is reduced to the constant and so (34) can be rewritten into
From (69), the assumption that and share IM, and the supposition that is a positive integer satisfying we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
and
are two entire functions. Next we use (20), (70), (71), Lemma 4, the lines of (44) and (45) in Subcase 1.2 to deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that one of and holds. Then, we use (69)-(71) and the similar reasonings as in Subcase 1.2 to get a contradiction. Next we suppose that and are non-vanishing polynomials. Then, we use (70), (71), the the similar reasonings from the line before (47) to the line of (51) to deduce
and
where and in what follows, and are two finite and non-zero complex constants. From (69) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with and for and Therefore, from (20), Lemma 2(ii), Remark 2, Lemma 4 and the first fundamental theorem, we deduce
and so we have
Since and are transcendental meromorphic functions such that and share CM, we deduce by (19), (74) and Lemma 5 that
From (72) and (73) we deduce that if is a common pole of and then is a common zero of and Combining this with (75) and the assumption that and share CM, we deduce
By (72), (73) and the Valiron-Mokhon′ko lemma (cf.[32]) we deduce
and
By (19), (76), (80), Lemma 5 and the obtained result that and are transcendental meromorphic functions we deduce
By (72), (73), (77), (81) and Lemma 7, we see that there exist some two integers and with such that This together with the definitions of and in (72) and (73) respectively gives
Next we use (82) and the similar reasonings from the line after (60) to the end of Subcase 1.2 to get a contradiction.
Case 2. Suppose that is a non-zero polynomial, then it follows by the definition of in (29) we see that is a non-constant polynomial. Next we let
where is a positive integer, and with are finite complex constants. Next we let with and and consider the following angles for any given positive number satisfying
where is an integer satisfying Then, it follows from (83) and Lemma 8 that there exists a positive number such that for we have
while
For convenience we next let
and
with
and
for each integer satisfying Then, we have the following claim:
Claim 2I. Based upon the assumptions of Theorem 7 and the supposition of Case 2, there are at most finitely many common zeros of and in for each integer and each integer satisfying
We prove Claim 2I: on the contrary, we suppose that Claim 2I is not valid. Then, there are infinitely many zeros of in for each integer and each integer satisfying Next we let be the infinite sequence of all the zeros of on for and the integer satisfying where each distinct point in is repeated as many times as its multiplicity of a zero of with and the infinite sequence of is arranged according to increasing moduli, such that
and
for We consider the following two subcases:
Subcase 2.1. Suppose that with and is an even integer. Then, by (30), (84), (90), (91) and (32) we deduce that there exist some positive integer and a large positive number satisfying for and the even integer satisfying such that
and such that each point satisfying (92) is not a common multiple zero of and with and satisfies
as and for and the even integer satisfying Here and in what follows, denotes the multiplicity of of a zero of for the positive integer satisfying and the even integer satisfying Therefore, from (30), (32)-(93) and the assumption that and share and IM, we deduce that is also a simple zero of for the positive integer satisfying and the even integer satisfying Next we use Lemma 10 for to deduce
for the positive number such that From (92), (93) and Lemma 9 we have
for the the even integer satisfying and the positive number such that From (20) and Lemma 12 we have
for the even integer satisfying and the positive number such that From (94)-(96) we deduce
for the even integer satisfying and the positive number such that Next we use Lemma 9 and the reasonings in the proof of Zheng [56, pp.53-54, Lemma 2.2.2] to deduce
for the even integer satisfying and the positive number such that Here and in what follows,
for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (97) and (98) we have
for and the positive number such that
Now we use (92) and (93) to deduce
for the positive integer satisfying Combining this with (90), we deduce that for any given positive number satisfying there exists some positive integer with that satisfies
such that for the positive integer satisfying with we have
From (102) we have
for the positive integer satisfying with From (100)-(103) we have
for the positive integer satisfying with From (90), (104) and (99) with the even integer satisfying we have
for the even integer satisfying and the given positive number satisfying This is a contradiction.
Subcase 2.2. Suppose that with and is an odd integer. Then, by (30), (85), (90), (91) and (32) we deduce that there exist some positive integer and a large positive number satisfying for and the odd integer satisfying such that (92) holds and each point satisfying (92) is not a common multiple zero of and for and each point satisfying (92) satisfies
as for and the odd integer satisfying Here and in what follows, denotes the multiplicity of of a zero of for the positive integer satisfying and the odd integer satisfying From (30), (32), (106) and the assumption that and share and IM, we deduce that is also a simple zero of for the positive integer satisfying and the odd integer satisfying Next we use Lemma 10 for to deduce
for and the odd integer satisfying From (92), (106) and Lemma 9 we have
for the the odd integer satisfying and the positive number such that From (20) and Lemma 12 we have
for the odd integer satisfying and the positive number such that From (107)-(109) we deduce
for the odd integer satisfying and the positive number such that Next we use Lemma 9 and the reasonings in the proof of Zheng [56, pp.53-54, Lemma 2.2.2] to have
for the odd integer satisfying and the positive number such that Here and in what follows, is defined as with being the number of zeros of for in counted with multiplicities. From (110) and (111) we have
for the odd integer satisfying and the positive number such that Now we use (92) and (106) to deduce for the positive integer satisfying Combining this with (90), we deduce that for the positive number satisfying there exists some large positive integer with that satisfies (101), such that for the positive integer satisfying with we have (102). From (102) we have
for and the positive integer satisfying From (101), (102), (112), (113) we have
for the odd integer satisfying and the positive integer satisfying with From (90), (114) and (99) with the odd integer satisfying we deduce (105) for the odd integer satisfying and the given positive number satisfying From (105) we get a contradiction. This proves Claim 2I.
Next we use Lemma 16, Lemma 18 and Claim 2I to complete the proof of Theorem 7. For this purpose, we have a discussion as follows: suppose that Then, it follows from Lemma 2.2 in [21] that there exist some real constant and some set that has infinite linear measure such that for all Combining this with Theorem 5, we see that and share all four values CM, and so Theorem 7 is valid. Next we suppose that Then, it follows from the known equality that
Next we define the following directional arc:
with
for each integer satisfying From (116) and (117) we have
On the other hand, from (20), (115) and Lemma 15 we have
From (115)-(119) we see that there exists some integer satisfying say such that
with
for arbitrary Pó1ya peaks of order as Here and in what follows, is some positive constant, and is some set of finite logarithmic measure. Then, it follows from (120) and (121) that
where and in what follows, is a large positive number satisfying
Next we set the simply connected domain
with the positive boundary where and are two oriented segments such that
and
respectively, and is a directional arc such that
where and in what follows, and are defined as (89) for Next we also define with as
for Here and in what follows, is a small positive number satisfying and
is a complex number such that Then, from (123)-(128), Remark 5.3, the positive boundary of the region and the definition of the proximity function of with the center at for we have
for the small positive number satisfying Next we also set the following transformation
where and in what follows, is a simply connected domain such that
where is a positive number satisfying
and the positive boundary of the simply connected domain is where and in what follows, is an oriented segment, and is a directional arc. They are defined as follows: the oriented segment is such an oriented segment that for each we have with and the starting point of is while the end point of is The directional arc is such a directional arc that for each we have with and the starting point of is with while the end point of is with In addition, from (89) with we have
From (130) and (133) we can easily verify that the transformation of (130) maps the simply connected domain of (131) onto the simply connected domain of (123) in a one-to-one mapping manner, and maps the positive boundary of into the positive boundary of in a one-to-one mapping manner, where the oriented segment and the directional arc are mapped into and respectively by the transformation of (130) in a one-to-one mapping manner. Moreover, from (131)-(133), the polar form of the Cauchy-Riemann partial differential equations and the polar form of the sufficient conditions for an analytic function (cf.[43]), we can verify that the transformation of (130) is also an analytic transformation in the simply connected domain of (131). Combining this with the known result that the transformation of (130) maps the simply connected domain of (131) onto the simply connected domain of (123) in a one-to-one mapping manner, we see that the univalent transformation (130) on the simply connected domain is a univalent and analytic map on the simply connected domain and maps conformally the simply connected domain of (131) with the positive boundary onto the simply connected domain of (123) with the positive boundary For convinience of the following discussion, we denote the univalent and analytic transformation (130) as
and there exists a unique point satisfying
such that the univalent and analytic mapping (134) maps conformally the simply connected domain of (131) with the positive boundary onto the simply connected domain of (123) with the positive boundary On the other hand, from the Riemann mapping theorem (cf.[2, p.230,Theorem 1]) we also see that there exists a unique univalent and analytic function on the simply connected domain of (123), say
such that the univalent and analytic transformation (136) maps conformally the simply connected domain of (123) with the positive boundary onto the open unit disk with the positive boundary with where the starting point of is with and the end point of is with From (134)-(136) we deduce that the composition
of the transformation (136) and the transformation (134) with for is a univalent and analytic transformation of onto the open unit disk such that it maps the positive boundary of into the positive boundary of the open unit disk mentioned above in a one-to-one mapping manner. Noting that the transformation of (136) is a univalent and analytic mapping of onto the open unit disk Combining this with the theorem on the boundary behavior of conformal mappings, we see that the function (137) is continuous on and for Next we use the similar reasoning of proving (8) in Remark 5 to have
and
for each point with and on the semi-circle
Since the transformation (134) is a univalent and analytic transformation of onto such that (135) holds, we have from (134), (136)-(138), (140) and Lemma 17 that the Green’s functions and for and with singularities and respectively satisfy the relation
for each Next we use (135), (137)-(142) and the invariance of first order differential form to deduce
with for each point with and on the semi-circle From (128), (131), (132), (141), (13) in Remark 5.1 and Goldberg-Ostrovskii [14, p.306] we have
for each point with and on the semi-circle and the point
Here and in what follows
From (122), (129), (135), (143), (144)-(146), the supposition the definitions of the proximity function and the Nevanlinna characteristic function of with the center at for for and we deduce
i.e.,
for the small positive number satisfying and the positive number and Here and in what follows,
Next we denote by the Nevanlinna′s angular characteristic function on the angular domain
with
Next we use (149), (150) and the line of the reasoning for proof of (97) in Subcase 2.1 of the proof of Claim 2I to deduce
for the positive number such that say. From (122), (149)-(151) and Lemma 13, we deduce that there exists some positive constant and there exists some set of finite logarithmic measure, such that when and the relation
holds uniformly in Here and in what follows
Similarly, from Lemma 13 and (97) for we deduce that there exists some positive constant and there exists some set of finite logarithmic measure, such that the relation
holds uniformly in when and Here and in what follows, is defined as (99) for From (128), (154) and the supposition we deduce
as and On the other hand, from (147), (150), (152) and (153) we have
as and Last, from Lemma 16 we have
where
with and for Noting that denotes the Green function for the half-disc
we use the formula of the Green function (12) for the half-disc
in Remark 5.1 to have
with
for each From (130), (142) and (159) we have
for each with From Claim 2I we see that there exist finitely many distinct zeros of with in say such that
where is some positive integer that depends only upon the positive numbers and Next we suppose that the multiplicity of of a zero of is equal to some positive integer for and Then, from (161), (146), the reasoning in Goldberg-Ostrovskii [14, p.306], the supposition and the definition (5) of the counting function for the meromorphic function with the center at for we have
as Here and in what follows,
for and On the other hand, from Lemma 18 we see that there exists some set of finite logarithmic measure such that
as and From the formula (2.1.1) in Zheng [56, p.27], we have
From (164), (165) and the definitions of the proximity functions and the Nevanlinna characteristic functions of and for with the center at for we have
as and From (128) and (155) we deduce
as and From Remark 4, Lemma 3 (i), the assumption and the assumption that and share IM and CM we have
which implies that and have finitely many critical values including and possibly in the complex plane. Combining this with Lemma 20 and the assumption we deduce that the set of finite singular values of the inverse function is bounded. Therefore, from Lemma 19 we see that there exist constants and such that if and then
where is a positive absolute constant, in particular independent of and From (128), (155), (168) and the supposition we see that there exists some set of finite logarithmic measure such that
as and From (133) and (99) for we have Combining this with (128), (155) and (169) we deduce
as and From (128), (155), (170), Lemma 18 and the supposition we see that there exists some set of finite logarithmic measure, such that
as and From (158), (166), (167), (171) we have
as and From (157), (158), (162), (163), (172) we have
as and as From (146), (156) and (173) we have
as and From (148), (153), (89) and (99) for we see that
for any given small positive number satisfying From (174), (175) and (146) we derive a contradiction, when and This completes the proof of Theorem 7.
4. On a question of Gary G. Gundersen concerning the nonexistence of two distinct and non-constant meromorphic functions sharing three distinct values DM and a fourth value CM
In 1979, G.G.Gundersen proposed the following two questions in Gundersen [15]:
Question 3.
([15, p.458]) Do there exist two distinct and non-constant meromorphic functions that share two distinct values CM and share the other two distinct values DM, where the four values are four distinct values in the extended complex plane?
Question 4.
([15, p.458]) Do there exist two distinct and non-constant meromorphic functions that share one value CM and share the other three distinct values DM, where the four values are four distinct values in the extended complex plane?
In 1983, Gundersen[16, Theorem 1] proved that if two distinct and non-constant meromorphic functions share two values CM and share two other values IM, where the four values are four distinct values in the extended complex plane, then the two meromorphic functions share all four values CM. This implies that Gundersen[16, Theorem 1] gave a negative answer to Question 3. But Question 4 is still open by now. The following question is a special case of Question 4:
Question 5.
Do there exist two distinct and non-constant entire functions that share three distinct finite values DM?
Theorem 8.
([34, pp.109-117]) There do not exist two distinct and non-constant entire functions that share three distinct finite values DM.
For the existence of three distinct and non-constant meromorphic functions sharing four distinct values IM, where the four distinct shared values are in the extended complex plane that are neither CM shared values nor DM shared values, we refer to Steinmetz [44, Theorem 2]. For the existence of two distinct and non-constant meromorphic functions sharing four distinct values DM, where the four distinct shared values are in the extended complex plane, we refer to Reinders [39, Theorem 1].
From Theorem 7 we get the following result that gives a negative answer to Question 4 for the finite order meromorphic functions:
Theorem 9.
There do not exist two distinct and non-constant finite order meromorphic functions that share DM and share CM, where are four distinct complex values in the extended complex plane.
From Example 1 above and the following example we see that the number of the three distinct DM shared values and the number of one CM shared value in Theorem 9, in a sense, are best possible:
Example 2.
([44, p.94]) Let denote the Weierstrass -function with a pair of primitive periods and Then
and share the value IM, where and are two distinct complex numbers such that with is a cube root of unity, while and are two complex numbers such that and are linearly independent on That is to say, for each pair of real numbers and such that we have Moreover, we can verify that the four values are neither CM shared values nor DM shared values of and
5. Concluding remarks
Conjecture 1.
Suppose that and are two distinct transcendental meromorphic functions of infinite order, if and share IM and CM, where are four distinct complex values in the extended complex plane, then and share the four values CM.
Conjecture 2.
There do not exist two distinct and non-constant infinite order meromorphic functions that share DM and share CM, where are four distinct complex values in the extended complex plane.